OpenStudy (anonymous):

Find the vertex of the graph of the function. f(x) = (x - 5)2 + 10

OpenStudy (anonymous):

\[f(x)=(x-5)^{2}+10\]

OpenStudy (madhu.mukherjee.946):

f(x) = a(x-h)^2 + k --->vertex = (h,k) so vertex is (5,10)

OpenStudy (anonymous):

(5,10)

OpenStudy (anonymous):

thank youuuuu

OpenStudy (anonymous):

what about this one? its a little different Find the vertex of the graph of the function. \[f(x) = 3x ^{2} - 18x + 24\]

OpenStudy (anonymous):

@madhu.mukherjee.946

OpenStudy (madhu.mukherjee.946):

f(x)=3x^2-18x+24 completing the square f(x)=3(x^2-6x+9)+24-27 f(x)=3(x-3)^2-3

OpenStudy (madhu.mukherjee.946):

This is a parabola of the standard form: A(x-h)^2+k, with (h,k) being the (x,y) coordinates of the vertex. Positive leading coefficient means parabola open upwards, has a minimum. For given parabola: vertex: (3,-3)

OpenStudy (anonymous):

do u need help still?

OpenStudy (anonymous):

thank you so much!! @madhu.mukherjee.946

OpenStudy (madhu.mukherjee.946):

your welcome darling:))

OpenStudy (anonymous):

no i think i got it but thank you @Skielerlucas04 ill tell you if i dont get something else

OpenStudy (anonymous):

ok